OpenStudy (anonymous):

Let a be a positive real number. Show that loga(x+√x^2−1)+loga(x−√x^2−1)=0

OpenStudy (misty1212):

combine as a single log using \[\log(A)+\log(B)=\log(AB)\] maybe if you then multiply it will be 1 insides, and \(\log(1)=0\)